# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v1

@@ 1 Introduction
Figure 1: Illustration of denoising equilibrium and non-equilibrium structures. In this figure, we relax a non-equilibrium structure (red point) and form a local relaxation trajectory (black dotted arrow). All the points along the trajectory except the blue point are non-equilibrium structures.
@@
To start addressing this gap, we take inspiration from self-supervised learning methods in NLP and CV and explore how we can adapt them to learn better atomistic representations from existing labeled data. Specifically, one of the most popular self-supervised learning methods in NLP (Devlin et al., 2019) and CV (He et al., 2022) is training a denoising autoencoder (Vincent et al., 2008), where the idea is to mask or corrupt a part of the input data and learn to reconstruct or denoise the original, uncorrupted data. Denoising assumes we know a unique target structure such as a sentence and an image in the case of NLP and CV. Indeed, this is the case for equilibrium structures (e.g., $S_{\text{eq}}$ at a local energy minimum in Figure 1b) as has been demonstrated by previous works leveraging denoising for pretraining on atomistic data (Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Wang et al., 2023; Feng et al., 2023a). However, most previous works are limited to equilibrium structures, and equilibrium structures constitute only a small portion of available data since structures along a trajectory to get to a local minimum are all non-equilibrium. Hence, it is important to generalize denoising to leverage the much larger set of non-equilibrium structures.
@@
Since a non-equilibrium structure has non-zero atomic forces and atoms are not confined to local energy minima, it can have more possible atomic positions than a structure at equilibrium. As shown in Figure 1a, this can make denoising a non-equilibrium structure an ill-posed problem since there are many possible target structures. To address the issue, we propose to take the forces of the original non-equilibrium structures as inputs when denoising non-equilibrium structures. Intuitively, the forces constrain the atomic positions of a non-equilibrium structure. With the additional information, we are able to predict the original non-equilibrium structure satisfying the input forces instead of predicting any arbitrary structures as shown in Figure 1c. Previous works on denoising equilibrium structures (Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; a) end up being a special case where the forces of original structures are close to zero.
@@
Based on the insight, in this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage atomistic data. For training DeNS, we first corrupt a structure by adding noise to its 3D atomic coordinates and then reconstruct the original uncorrupted structure by predicting the noise. For noise predictions, a model is given the forces of the original uncorrupted structure as inputs to make the transformation from a corrupted non-equilibrium structure to an uncorrupted non-equilibrium structure tractable. When used along with the original tasks like predicting the energy and forces of non-equilibrium structures, DeNS improves the performance of the original tasks with a marginal increase in training cost. We further discuss how DeNS can better leverage training data to improve performance and show the connection to self-supervised learning methods in other domains.
@@
We mainly focus on how DeNS can improve equivariant networks and conduct extensive experiments on OC20 (Chanussot* et al., 2021), OC22 (Tran* et al., 2022) and MD17 (Chmiela et al., 2017; Schütt et al., 2017; Chmiela et al., 2018) datasets. EquiformerV2 (Liao et al., 2023) trained with DeNS achieves better S2EF results and comparable IS2RE results on OC20 (Section 4.1). EquiformerV2 trained with DeNS sets new state-of-the-art results on OC22, improving energy by ${\sim}13\%$, forces by ${\sim}7\%$, and IS2RE by ${\sim}12\%$ (Section 4.2). On MD17, Equiformer ($L_{max}=2$(Liao & Smidt, 2023) trained with DeNS achieves better results and saves 3.1×3.1$\times$3.1 × training time compared to Equiformer ($L_{max}=3$) without DeNS (Section 4.3), where $L_{max}$ denotes the maximum degree. Additionally, DeNS can improve other equivariant networks like eSCN (Passaro & Zitnick, 2023) on OC20 and SEGNN-like networks (Brandstetter et al., 2022) on MD17.
@@ 2 Related Works
Denoising structures have been used to boost the performance of GNNs on 3D atomistic datasets (Godwin et al., 2022; Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; Wang et al., 2023; Feng et al., 2023a). The approach is to first corrupt data by adding noise and then train a denoising autoencoder to reconstruct the original data by predicting the noise, and the motivation is that learning to reconstruct data enables learning generalizable representations (Devlin et al., 2019; He et al., 2022; Godwin et al., 2022; Zaidi et al., 2023). Since denoising equilibrium structures do not require labels and is self-supervised, similar to BERT (Devlin et al., 2019) and MAE (He et al., 2022), it is common to pre-train via denoising on a large dataset of equilibrium structures like PCQM4Mv2 (Nakata & Shimazaki, 2017) and then fine-tune with supervised learning on smaller downstream datasets. Besides, Noisy Nodes (Godwin et al., 2022) use denoising equilibrium structures as an auxiliary task along with original tasks without pre-training on another larger dataset. However, most of the previous works are limited to equilibrium structures, which occupy a much smaller amount of data than non-equilibrium ones. In contrast, the proposed DeNS generalizes denoising to non-equilibrium structures with force encoding so that we can improve the performance on the larger set of non-equilibrium structures. We provide a detailed comparison to previous works on denoising in Section A.2.
@@
Refer to Section A.1 for discussion on equivariant networks.
@@ 3.1 Problem Setup
where $\lambda_{E}$ and $\lambda_{F}$ are energy and force coefficients controlling the relative importance between energy and force predictions. $E^{\prime}(S_{\text{non-eq}})=\frac{E(S_{\text{non-eq}})-\mu_{E}}{\sigma_{E}}$ is the normalized ground truth energy obtained by first subtracting the original energy $E(S_{\text{non-eq}})$ by the mean of energy labels in the training set $\mu_{E}$ and then dividing by the standard deviation of energy labels $\sigma_{E}$. Similarly, $\mathbf{f}_{i}^{\prime}=\frac{\mathbf{f}_{i}}{\sigma_{F}}$ is the normalized atom-wise force. For force prediction, we can either directly predict them from latent representations like node embeddings as commonly used for OC20 and OC22 datasets or take the negative gradients of predicted energy with respect to atomic positions for datasets like MD17.
@@ 3.2.1 Formulation of Denoising
Denoising structures have been used to improve the performance of GNNs on 3D atomistic datasets. They first corrupt data by adding noise and then train a denoising autoencoder to reconstruct the original data by predicting the noise. Specifically, given a 3D atomistic system $S=\left\{(z_{i},\mathbf{p}_{i})\mid i\in\{1,...,|S|\}\right\}$, we create a corrupted structure $\tilde{S}$ by adding Gaussian noise with a tuneable standard deviation $\sigma$ to the atomic positions $\mathbf{p}_{i}$ of the original structure $S$:
@@
We denote the set of noise added to $S$ as $\text{Noise}(S,\tilde{S})=\left\{\bm{\epsilon_{i}}\in\mathbb{R}^{3}\mid i\in\{1,...,|S|\}\right\}$. When training a denoising autoencoder, GNNs take $\tilde{S}$ as inputs, output atom-wise noise predictions $\hat{\bm{\epsilon}}(\tilde{S})_{i}$ and minimize the L2 difference between normalized noise $\frac{\bm{\epsilon}_{i}}{\sigma}$ and noise predictions $\hat{\bm{\epsilon}}(\tilde{S})_{i}$:
@@
where $p(S,\tilde{S})$ denotes the probability of obtaining the corrupted structure $\tilde{S}$ from the original structure $S$. We divide the noise $\bm{\epsilon}_{i}$ by the standard deviation $\sigma$ to normalize the outputs of noise predictions.
When the original structure $S$ is an equilibrium structure, denoising is to find the structure corresponding to the nearest energy local minima given a high-energy corrupted structure. This makes denoising equilibrium structures a many-to-one mapping and therefore a well-defined problem. However, when the original structure $S$ is a non-equilibrium structure, denoising, the transformation from a corrupted non-equilibrium structure to the original non-equilibrium one, can be an ill-posed problem since there are many possible target structures as shown in Figure 1a.
@@ 3.2.2 Force Encoding
The reason that denoising non-equilibrium structures can be ill-posed is because we do not provide sufficient information to specify the properties of the target structures. Concretely, given an original non-equilibrium structure $S_{\text{non-eq}}$ and its corrupted counterpart $\tilde{S}_{\text{non-eq}}$, some structures interpolated between $S_{\text{non-eq}}$ and $\tilde{S}_{\text{non-eq}}$ could be in the same data distribution and therefore be the potential target structures of denoising. In contrast, when denoising equilibrium structures as shown in Figure 1b, we implicitly provide the extra information that the target structure should be at equilibrium with near-zero forces, and this therefore limits the possibility of the target of denoising. Motivated by the assumption that the forces of the original structures are close to zeros when denoising equilibrium ones, we propose to encode the forces of original non-equilibrium structures when denoising non-equilibrium ones as illustrated in Figure 1c. Specifically, when training denoising non-equilibrium structures (DeNS), GNNs take both a corrupted non-equilibrium structure $\tilde{S}_{\text{non-eq}}$ and the forces $F(S_{\text{non-eq}})$ of the original non-equilibrium structure $S_{\text{non-eq}}$ as inputs, output atom-wise noise predictions $\hat{\bm{\epsilon}}\left(\tilde{S}_{\text{non-eq}},F(S_{\text{non-eq}})\right)_{i}$ and minimize the L2 difference between normalized noise $\frac{\bm{\epsilon}_{i}}{\sigma}$ and noise predictions $\hat{\bm{\epsilon}}\left(\tilde{S}_{\text{non-eq}},F(S_{\text{non-eq}})\right)_{i}$:
@@
Equation 4 is more general and reduces to Equation 3 when the target of denoising becomes equilibrium structures with near-zero forces. Since we train GNNs with $\tilde{S}_{\text{non-eq}}$ and $F(S_{\text{non-eq}})$ as inputs and $\text{Noise}(S_{\text{non-eq}},\tilde{S}_{\text{non-eq}})$ as outputs, they implicitly learn to leverage $F(S_{\text{non-eq}})$ to reconstruct $S_{\text{non-eq}}$ instead of predicting any arbitrary non-equilibrium structures. Empirically, we find that force encoding is critical to the effectiveness of DeNS on OC20 S2EF-2M dataset (Index 1 and Index 2 in Table LABEL:tab:design_choices) and MD17 dataset (Index 2 and Index 3 in Table 6).
@@
Figure 2: Training process when incorporating DeNS as an auxiliary task. The upper blue block corresponds to the original task of energy and force predictions (Equation 1), and the lower red block corresponds to training DeNS (Equation 6). “Equivariant GNN” and “energy head” are shared across the two tasks. For each batch of structures, we use the original task for some structures and DeNS for the others.
@@ 3.2.3 Training with DeNS
We propose to use DeNS as an auxiliary task along with the original task of predicting energy and forces and summarize the training process in Figure 2. Specifically, given a batch of structures, for each structure, we decide whether we optimize the objective of DeNS (Equation 6) or the objective of the original task (Equation 1). This introduces an additional hyper-parameter $p_{\text{DeNS}}$, the probability of optimizing DeNS. We use an additional noise head for noise prediction, which slightly increases training time. Additionally, when training DeNS, similar to Noisy Nodes (Godwin et al., 2022), we also leverage energy labels and predict the energy of original structures. Therefore, given an original non-equilibrium structure $S_{\text{non-eq}}$ and the corrupted counterpart $\tilde{S}_{\text{non-eq}}$, training DeNS corresponds to minimizing the following loss function:
@@
where $\mathcal{L}_{\text{DeNS}}$ denotes the loss function of denoising as defined in Equation 4. We note that we also encode forces as discussed in Section 3.2.2 to predict the energy of $S_{\text{non-eq}}$, and we share the energy prediction head across Equation 1 and Equation 6. The loss function introduces another hyper-parameter $\lambda_{\text{DeNS}}$, DeNS coefficient, controlling the relative importance of the auxiliary task. Besides, the process of corrupting structures also results in another hyper-parameter $\sigma$ as shown in Equation 2. We provide the pseudocode in Section E. We note that we only train DeNS with force encoding on the training set without using any force labels on the validation and testing sets.
@@
Empirically, we find that adding noise to all atoms in a structure can lead to limited performance gain of DeNS on OC22 and MD17 datasets, and we surmise that there are still several structures satisfying input forces if we add noise to all atoms, making denoising too difficult a task and potentially not well-defined. To address this issue, we use partially corrupted structures, where we only add noise to and denoise a subset of atoms. Specifically, for corrupted atoms, we encode their atom-wise forces and predict the noise. For other uncorrupted atoms, we do not encode forces and train for the original task of predicting forces. We also predict the energy of the original structures given partially corrupted structures. This introduces an additional hyper-parameter, corruption ratio $r_{\text{DeNS}}$, which is the ratio of the number of corrupted atoms to that of all atoms.
Multi-Scale Noise.
Inspired by prior works on denoising score matching (Song & Ermon, 2019; 2020), we empirically find that incorporating multiple noise scales together for denoising improves energy and force predictions on OC20 and OC22 datasets. Specifically, we choose the standard deviations $\{\sigma_{k}\}_{k=1}^{T}$ to be a geometric progression that satisfy $\frac{\sigma_{T}}{\sigma_{T-1}}=...=\frac{\sigma_{3}}{\sigma_{2}}=\frac{\sigma_{2}}{\sigma_{1}}>1$:
$\sigma_{k}=\text{exp}\left(\log_{e}\sigma_{\text{low}}+\frac{k-1}{T-1}\cdot\left(\log_{e}\sigma_{\text{high}}-\log_{e}\sigma_{\text{low}}\right)\right)\quad\text{where}\quad k=1,...,T$ (7)
Here we use $\sigma_{1}=\sigma_{\text{low}}=0.01$ and $T=50$ and only tune $\sigma_{T}=\sigma_{\text{high}}$ when multi-scale noise is adopted. When training with DeNS, for each structure, we first sample a single noise standard deviation $\sigma$ from the $T$ values, and then follow Equations 2 and 4. We surmise that multi-scale noises are more likely to span the distribution of meaningful non-equilibrium structures across a diverse range of geometries compared to a fixed noise standard deviation.

### v2

@@ 1 Introduction
Figure 1: Illustration of denoising equilibrium and non-equilibrium structures. As shown in (a), we relax a non-equilibrium structure $S_{\text{non-eq}}$ and obtain the final relaxed, equilibrium structure $S_{\text{eq}}$. The path between $S_{\text{non-eq}}$ and $S_{\text{eq}}$ forms a relaxation trajectory. All structures $S$ along the trajectory except $S_{\text{eq}}$ are non-equilibrium and have non-zero forces $F(S)$. For denoising structures $S$, we add noise to their 3D atomic coordinates to obtain corrupted structures $\tilde{S}$ and predict the corresponding noise given $\tilde{S}$. We compare denoising non-equilibrium and equilibrium structures in (b) and (c), respectively, and show that denoising non-equilibrium structures can be ill-posed in (b). The issue in (b) can be addressed with force encoding, where we take forces as additional inputs, as in (d) and (e).
@@
To start addressing this gap, we take inspiration from self-supervised learning methods in NLP and CV and explore how we can adapt them to learn better atomistic representations from existing labeled data. Specifically, one of the most popular self-supervised learning methods in NLP (Devlin et al., 2019) and CV (He et al., 2022) is training a denoising autoencoder (Vincent et al., 2008), where the idea is to mask or corrupt a part of the input data and learn to reconstruct or denoise the original, uncorrupted data. Denoising assumes we know a unique target structure such as a sentence and an image in the case of NLP and CV. Indeed, this is the case for equilibrium structures (e.g., $S_{\text{eq}}$ at a local energy minimum in Figure 1(c)) as has been demonstrated by previous works leveraging denoising for pretraining on atomistic data (Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Wang et al., 2023; Feng et al., 2023a). However, most previous works are limited to equilibrium structures, and equilibrium structures constitute only a small portion of available data since structures along a relaxation trajectory to get to a local minimum are all non-equilibrium as shown in Figure 1. Hence, it is important to generalize denoising to leverage the much larger set of non-equilibrium structures.
@@
Since a non-equilibrium structure has non-zero atomic forces and atoms are not confined to local energy minima, it can have more possible atomic positions than a structure at equilibrium. As shown in Figure 1(b), this can make denoising a non-equilibrium structure an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of the original non-equilibrium structures as inputs when denoising non-equilibrium structures. Intuitively, the forces constrain the atomic positions of a non-equilibrium structure. With the additional information, we are able to predict the original non-equilibrium structure satisfying the input forces instead of predicting any arbitrary structures as shown in Figure 1(d). Previous works on denoising equilibrium structures (Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; a) end up being a special case where the forces of original structures are close to zero as in Figure 1(e).
@@
Based on the insight, in this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage atomistic data. For training DeNS, we first corrupt a structure by adding noise to its 3D atomic coordinates and then reconstruct the original uncorrupted structure by predicting the noise. For noise prediction, a model is given the forces of the original uncorrupted structure as inputs to make the transformation from a corrupted non-equilibrium structure to an uncorrupted non-equilibrium structure tractable. When used along with the original tasks like predicting the energy and forces of non-equilibrium structures, DeNS improves the performance on the original tasks with a marginal increase in training cost. We further discuss how DeNS can better leverage training data to improve performance and show the connection to self-supervised learning methods in other domains.
@@
We mainly focus on how DeNS can improve equivariant networks and conduct extensive experiments on OC20 (Chanussot* et al., 2021), OC22 (Tran* et al., 2022) and MD17 (Chmiela et al., 2017; Schütt et al., 2017; Chmiela et al., 2018) datasets. On OC20 S2EF-2M dataset, EquiformerV2 (Liao et al., 2023) trained with DeNS can achieve better energy and force results and save 2.3×2.3$\times$2.3 × training time compared to training without DeNS (Section 4.1.1). On OC20 S2EF-All+MD dataset, EquiformerV2 trained with DeNS achieves new state-of-the-art results on Structure to Energy and Forces (S2EF) and Initial Structure to Relaxed Energy (IS2RE) tasks (Section 4.1.2). Similarly, EquiformerV2 trained with DeNS sets new state-of-the-art results on OC22 dataset, improving energy by up to $15\%$, forces by up to $12\%$, and IS2RE by up to $15\%$ (Section 4.2). On MD17 dataset, Equiformer ($L_{max}=2$(Liao & Smidt, 2023) trained with DeNS achieves better results and saves 3.1×3.1$\times$3.1 × training time compared to Equiformer ($L_{max}=3$) without DeNS (Section 4.3), where $L_{max}$ denotes the maximum degree. Additionally, DeNS can improve other equivariant networks like eSCN (Passaro & Zitnick, 2023) on OC20 and SEGNN-like networks (Brandstetter et al., 2022) on MD17.
@@ 2 Related Works
Denoising structures have been used to boost the performance of GNNs on 3D atomistic datasets (Godwin et al., 2022; Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; Wang et al., 2023; Feng et al., 2023a). The approach is to first corrupt data by adding noise and then train a denoising autoencoder to reconstruct the original data by predicting the noise, and the motivation is that learning to reconstruct data enables learning generalizable representations (Devlin et al., 2019; He et al., 2022; Godwin et al., 2022; Zaidi et al., 2023). Since denoising equilibrium structures do not require labels and is self-supervised, similar to BERT (Devlin et al., 2019) and MAE (He et al., 2022), it is common to pre-train via denoising on a large dataset of equilibrium structures like PCQM4Mv2 (Nakata & Shimazaki, 2017) and then fine-tune with supervised learning on smaller downstream datasets. Besides, the work of Noisy Nodes (Godwin et al., 2022) uses denoising equilibrium structures as an auxiliary task along with original tasks without pre-training on another larger dataset. However, most of the previous works are limited to equilibrium structures, which occupy a much smaller amount of data than non-equilibrium ones. In contrast, the proposed DeNS generalizes denoising to non-equilibrium structures with force encoding so that we can improve the performance on the larger set of non-equilibrium structures. We provide a detailed comparison to previous works on denoising in Section A.1.
@@
Please refer to Section A.2 for discussion on equivariant networks.
@@ 3.1 Problem Setup
where $\lambda_{E}$ and $\lambda_{F}$ are energy and force coefficients controlling the relative importance between energy and force predictions. $E^{\prime}(S_{\text{non-eq}})=\frac{E(S_{\text{non-eq}})-\mu_{E}}{\sigma_{E}}$ is the normalized ground truth energy obtained by first subtracting the original energy $E(S_{\text{non-eq}})$ by the mean of energy labels in the training set $\mu_{E}$ and then dividing by the standard deviation of energy labels $\sigma_{E}$. Similarly, $\mathbf{f}_{i}^{\prime}=\frac{\mathbf{f}_{i}}{\sigma_{F}}$ is the normalized atom-wise force. For force prediction, we can either use direct methods, which is to directly predict forces from latent representations like node embeddings, as commonly used for OC20 and OC22 datasets or gradient methods, which is to take the negative gradients of predicted energy with respect to atomic positions, for datasets like MD17.
@@ 3.2.1 Formulation of Denoising
Denoising structures has been used to improve the performance of GNNs on 3D atomistic datasets (Godwin et al., 2022; Zaidi et al., 2023; Feng et al., 2023b; Wang et al., 2023). We first corrupt data by adding noise and then train a denoising autoencoder to reconstruct the original data by predicting the noise. Specifically, given a 3D atomistic system $S=\left\{(z_{i},\mathbf{p}_{i})\mid i\in\{1,...,|S|\}\right\}$, we create a corrupted structure $\tilde{S}$ by adding Gaussian noise with a tuneable standard deviation $\sigma$ to the atomic positions $\mathbf{p}_{i}$ of the original structure $S$:
@@
We denote the set of noise added to $S$ as $\text{Noise}(S,\tilde{S})=\left\{\bm{\epsilon_{i}}\in\mathbb{R}^{3}\mid i\in\{1,...,|S|\}\right\}$. When training a denoising autoencoder, GNNs take $\tilde{S}$ as inputs, output atom-wise noise prediction $\hat{\bm{\epsilon}}(\tilde{S})_{i}$ and minimize the L2 difference between normalized noise $\frac{\bm{\epsilon}_{i}}{\sigma}$ and noise prediction $\hat{\bm{\epsilon}}(\tilde{S})_{i}$:
@@
where $p(S,\tilde{S})$ denotes the probability of obtaining the corrupted structure $\tilde{S}$ from the original structure $S$. We divide the noise $\bm{\epsilon}_{i}$ by the standard deviation $\sigma$ to normalize the outputs of noise prediction.
When the original structure $S$ is an equilibrium structure, denoising is to find the structure corresponding to the nearest energy local minima given a high-energy corrupted structure. This makes denoising equilibrium structures a many-to-one mapping and therefore a well-defined problem. However, when the original structure $S$ is a non-equilibrium structure, denoising, the transformation from a corrupted non-equilibrium structure to the original non-equilibrium one, can be an ill-posed problem since there are many possible target structures as shown in Figure 1(b).
@@ 3.2.2 Force Encoding
The reason that denoising non-equilibrium structures can be ill-posed is because we do not provide sufficient information to specify the properties of the target structures. Concretely, given an original non-equilibrium structure $S_{\text{non-eq}}$ and its corrupted counterpart $\tilde{S}_{\text{non-eq}}$, some structures interpolated between $S_{\text{non-eq}}$ and $\tilde{S}_{\text{non-eq}}$ could be in the same data distribution and therefore be the potential target structures of denoising. In contrast, when denoising equilibrium structures as shown in Figure 1(c), we implicitly provide the extra information that the target structure should be at equilibrium with near-zero forces, and this therefore limits the possibility of the target of denoising. Motivated by the assumption that the forces of the original structures are close to zeros when denoising equilibrium ones, we propose to encode the forces of original non-equilibrium structures when denoising non-equilibrium ones as illustrated in Figure 1(d). Specifically, when training denoising non-equilibrium structures (DeNS), GNNs take both a corrupted non-equilibrium structure $\tilde{S}_{\text{non-eq}}$ and the forces $F(S_{\text{non-eq}})$ of the original non-equilibrium structure $S_{\text{non-eq}}$ as inputs, output atom-wise noise prediction $\hat{\bm{\epsilon}}\left(\tilde{S}_{\text{non-eq}},F(S_{\text{non-eq}})\right)_{i}$ and minimize the L2 difference between normalized noise $\frac{\bm{\epsilon}_{i}}{\sigma}$ and noise prediction $\hat{\bm{\epsilon}}\left(\tilde{S}_{\text{non-eq}},F(S_{\text{non-eq}})\right)_{i}$:
@@
Equation 4 is more general and reduces to Equation 3 when the target of denoising becomes equilibrium structures with near-zero forces. Since we train GNNs with $\tilde{S}_{\text{non-eq}}$ and $F(S_{\text{non-eq}})$ as inputs and $\text{Noise}(S_{\text{non-eq}},\tilde{S}_{\text{non-eq}})$ as outputs, they implicitly learn to leverage $F(S_{\text{non-eq}})$ to reconstruct $S_{\text{non-eq}}$ instead of predicting any arbitrary non-equilibrium structures. Empirically, we find that force encoding is critical to the effectiveness of DeNS on OC20 S2EF-2M dataset (Index 2 and Index 3 in Table 1(b)) and MD17 dataset (Index 2 and Index 3 in Table 6).
@@
(a) Original task of predicting energy and forces.
(b) DeNS without partially corrupted structures. All the atoms in a structure are corrupted with Gaussian noise. We encode all the atom-wise forces to predict noise.
(c) DeNS with partially corrupted structures. We add noise to and denoise a random subset of atoms. For corrupted atoms, we encode their forces and predict the noise. For other uncorrupted atoms, we instead predict their forces. Here only the top two white atoms are corrupted.
Figure 2: Training process when incorporating DeNS as an auxiliary task. The pseudocode for training with DeNS can be found in Section E. “Equivariant GNN”, “energy head”, “force head” and “noise head” are shared across (a), (b) and (c). For each batch of structures, we use the original task (a) for some structures and DeNS ((b) or (c)) for the others. Using partially corrupted structures as in (c) is empirically better than (b). We note that the force label $F(S_{\text{non-eq}})$ and energy label $E(S_{\text{non-eq}})$ used in (b) and (c) are the same as those in (a) and that training with DeNS does not require any additional data.
@@ 3.2.3 Training with DeNS
We propose to use DeNS as an auxiliary task along with the original task of predicting energy and forces and summarize the training process in Figure 2. Specifically, given a batch of structures, for each structure, we decide whether we optimize the objective of DeNS (Figure 2(b) or Figure 2(c)) or the objective of the original task (Figure 2(a)). This introduces an additional hyper-parameter $p_{\text{DeNS}}$, the probability of optimizing DeNS. We use an additional noise head for noise prediction, which slightly increases training time. Additionally, when training DeNS, similar to Noisy Nodes (Godwin et al., 2022), we also leverage energy labels and predict the energy of original structures. Therefore, given an original non-equilibrium structure $S_{\text{non-eq}}$ and the corrupted counterpart $\tilde{S}_{\text{non-eq}}$, training DeNS corresponds to minimizing the following loss function:
@@
where $\mathcal{L}_{\text{DeNS}}$ denotes the loss function of denoising as defined in Equation 4. We note that we also encode forces as discussed in Section 3.2.2 to predict the energy of $S_{\text{non-eq}}$, and we share the energy prediction head across Equation 1 and Equation 6. The loss function introduces another hyper-parameter $\lambda_{\text{DeNS}}$, DeNS coefficient, controlling the relative importance of the auxiliary task. Besides, the process of corrupting structures also results in another hyper-parameter $\sigma$ as shown in Equation 2. We provide the pseudocode in Section E. We note that the force encoding in DeNS only relies on force labels on the training set and we do not use any force labels on the validation or testing sets.
@@
Empirically, we find that adding noise to all atoms in a structure can sometimes lead to limited performance gain of DeNS. We surmise adding noise to all atoms makes denoising too difficult and potentially not well-defined and that there are still several structures satisfying input forces. To address this issue, we use partially corrupted structures, where we only add noise to and denoise a random subset of atoms as shown in Figure 2(c). Specifically, for corrupted atoms, we encode their atom-wise forces and predict the noise. For other uncorrupted atoms, we do not encode forces and train for the original task of predicting forces. We also predict the energy of the original structures given partially corrupted structures. This introduces an additional hyper-parameter, corruption ratio $r_{\text{DeNS}}$, which is the ratio of the number of corrupted atoms to that of all atoms.